OpenStudy (bekkah323):

One plane travels 80 mph faster than another. while one travels 1750 mi, the other travels 950 mi. Find the speed of each plane.

OpenStudy (goformit100):

@hartnn I think it's a Physics Question... it's like a Spam here on Maths section

OpenStudy (goformit100):

@AccessDenied

OpenStudy (bekkah323):

no its a math question. I'm supposed to use the d=rt formula.

OpenStudy (whpalmer4):

\[d = r t\] Plane 1 has a speed of \(r_1\), Plane 2 has a speed of \(r_2 = r_1+80\) \[d_1 = r_1 t = 950\]\[d_2 = r_2 t = (r1+80) t = r_1 t + 80 t = 1750\]Substitute first equation in second:\[r_1 t + 80 t = 1750\]\[950 + 80t = 1750\]Solve for \(t\), and use \(t\) in both equations to find \(r_1, r_2\)

OpenStudy (whpalmer4):

if this isn't suitable for the math section, neither are all of the coin problems, interest problems, exponential decay/growth problems, etc. no particularly deep knowledge of physics required here — any math student should be able to do this problem.